OpenStudy (anonymous):

If sinΘ=4/5, what is the value of cosΘ in quadrant 1?

jimthompson5910 (jim_thompson5910):

Hint: \[\Large \sin^2(\theta)+\cos^2(\theta) = 1\]

jimthompson5910 (jim_thompson5910):

Since the terminal angle theta is in quadrant I, this means that \(\Large \cos(\theta) > 0\)

OpenStudy (mathmale):

I agree with Jim! Another way to solve this problem would be to draw the triangle in question. You'd need to use the Pyth. Thm. to find the length of the adjacent side; from there it's easy to write the cosine of theta. Because the sine of theta is 4/5, we already know that the opp side is 4 and the hyp is 5.

OpenStudy (mathmale):

If you go that route, you would, of course, want to draw the triangle in Quadrant 1.